An object thrown vertically up from the ground passes the height 5m twice in an interval of 10sec . What is the time of flight

Grade:11

1 Answers

Manas Choudhary
8 years ago
after reaching h=5m for the first time, the body travelled for 5 more secs.
v=u+at,u=50m/s *u is its velocity at h=5m.
considering the whole scenario let w be the velocity supplied to it.
solving further i got T=time of flight =1041/2 =2root 26=10.2 sec aproxx.